What would be the square root of 625 using the identity (a +b)2 = a2 + b2 + 2ab?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

(625)^2

= (600 + 25)^2

= 600^2+ 2 x 600 x 25 +25^2

= 360000 + 30000 + 625

= 390625


Thanks